# Img2CAD: Reverse Engineering 3D CAD Models from Images through VLM-Assisted Conditional Factorization

### v1

@@ 7. Applications
7.1. Image To CAD in the Wild
Although our model is trained exclusively on synthetic ShapeNet CAD-ified models, it demonstrates the ability to generalize to real-world images. This is achieved through the foundational knowledge provided by vision language models such as GPT-4V and CLIP. Figures 11 and 12 present qualitative results using images from Pix3D (Sun et al., 2018) and Google product images, which offer an extensive collection of indoor furniture. It is important to note that this task is exceptionally challenging due to the difficulties associated with generalization and CAD representation complexity.
7.2. Structure-Aware Shape Editing in the Wild
A significant advantage of our method is its capability for structure-aware editing in real-world scenarios, enabled by our CAD representation. Figure 13 illustrates how users can easily modify the CAD structure of an object in an image, guided by natural language instructions. The process involves first obtaining the CAD program corresponding to the input image. Then, we feed this CAD program back to GPT, asking it to identify the key modifications needed to achieve the desired outcome. Users can make structural changes, such as adding an arc, or continuous attribute changes, such as editing the leg length, thereby creating variations easily from an input image. This demonstrates the practical utility of our method for intuitive and precise CAD model editing.
@@ 8. Conclusion
We propose a novel approach for reverse engineering 3D CAD models from input images by conditionally factorizing this challenging task into two sub-problems. First, we leverage large foundation models (i.e., GPT-4V) to infer the discrete global structure with semantic information. We then introduce TrAssembler, which, conditioned on this discrete global structure, predicts the continuous attribute values of the CAD program. Additionally, we constructed a CAD dataset of common objects from ShapeNet to support the training of the continuous attribute prediction network. Our experiments demonstrate promising results, significantly outperforming previous baselines. Our method also makes a significant initial step towards generating CAD models from images in the wild, showcasing its potential for real-world applications.
@@
Acknowledgements. This work is supported by ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a Apple Scholars in AI/ML PhD Fellowship. Yang You is also supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University.
Appendix
A.1. Experiment Details
A.1.1. Sketch and Extrusion Command Details
In our experiments, to ensure compatibility with the OpenCasCade (Capgemini, [n. d.]) library, we always append the $OneSided$ indicator after the extrusion command, as we only consider one-sided extrusion. Additionally, there is a clockwise/counter-clockwise flag for the $arc$ command, which we always set to 1, as our sketches are always drawn in a counter-clockwise order.
A.1.2. GPT Prompt
We use the following text as the inference prompt for GPT-4V. For clarity, we also add explanations of the role of each section of the prompt as comments, all marked in red (these comments are not part of the prompt). In the comments, we also specify the paragraphs that correspond to Reminder, ContExample, and CoT used in our ablation studies in Table 4.
{mdframed}
[backgroundcolor=gray!20!white,roundcorner=10pt]# Basic rules of a sketch-extrude CAD program
I provide a new type of language to create CAD models from commands. The command has two types: sketch and extrude.
Sketch commands are used to specify closed curves on a 2D plane in 3D space. Primarily, we consider closed curves which are formed from concatenations of LINES, ARCS, and CIRCLES.
In our representation, a PROFILE is a list which starts with an indicator command $<SOL>$, followed by a series of curve commands $C_{i}$ (which are lines, arcs, or circles). There are four important requirements in creating our sketch profile. 1) The first curve starts at the ORIGIN. 2) The last curve ends at the ORIGIN. 3) The end of each curve is the start of the next curve. 4) The sketch should be drawn in COUNTERCLOCKWISE order.
Each curve command $C_{i}$ is described by its curve type $t_{i}\in{<SOL>,L,A,R}$, where $L$ is specified in $(x,y)$ format giving the next line end-point, $A$ is in $(x,y,\alpha,f)$ giving the next arc end-point, sweep angle $\alpha$ and counter-clockwise flag $f$ (0 clockwise, 1 counter-clockwise), $R$ is in $(x,y,r)$ representing the center $(x,y)$ and the radius $r$.
Extrude commands have two purposes. 1) It extrudes a sketch profile from a 2D plane into a 3D body with one-sided extrusion. 2) The command also specifies how to merge newly extruded 3D body with the previously created shape, using NewBody. Extrude commands are in the following format: $(n_{x},n_{y},n_{z},x,y,z,e,b,u)$, where $n_{x},n_{y},n_{z}$ represent the 2D plane’s orientation (normal) in 3D space. $x,y,z$ are the location of the plane’s origin, $e$ is the one-sided extrusion distance. The extrusion distance gives the HEIGHT or THICKNESS of the extruded profile ALONG THE EXTRUSION AXIS DIRECTION. $b$ is the boolean type which should be NewBody, $u$ is the extrusion type which should be OneSided.
It is also important to note the COORDINATE SYSTEM CONVENTION of the program outputs. The COORDINATE SYSTEM CONVENTION is given as the Z-AXIS is the gravity axis pointing DOWNWARDS, the X-AXIS points OUTWARDS, and the Y-AXIS points TO THE RIGHT.
# A CAD program example on cubes
For example, to create three cubes of side length 2 units placed horizontally in the x direction (with a space of 3 units between corresponding corners of adjacent cubes), we could use the following program:
¡SOL¿
L: (2,0)
L: (2,2)
L: (0,2)
L: (0,0)
E: (0,0,1,0,0,0,2,NewBody,OneSided)
¡SOL¿
L: (2,0)
L: (2,2)
L: (0,2)
L: (0,0)
E: (0,0,1,3,0,0,2,NewBody,OneSided)
¡SOL¿
L: (2,0)
L: (2,2)
L: (0,2)
L: (0,0)
E: (0,0,1,6,0,0,2,NewBody,OneSided)
Note how here, we also make sure to draw the sketches in COUNTERCLOCKWISE order. Moreover we see that the cubes are extruded along the z-axis,the gravity axis.
Since cuts are not covered by the above commands, there is a separate syntax to form cuts. If you want to cut out of a base sketch, you should insert ¡CUT¿ followed by the sketch (in the same syntax as above for usual sketches) you wish to cut out of the base sketch. The coordinates will be in the same frame. YOU MUST PLACE THIS CUT BLOCK AFTER THE SKETCH COMMANDS FOR THE BASE SKETCH, BUT BEFORE THE EXTRUDE COMMAND FOR THE BASE SKETCH.
IT’S ALSO IMPORTANT TO RECOGNIZE CURVES WHEN THEY ARISE IN SKETCHES, AND CREATE THEM WITH ARC COMMANDS. Arc commands can be useful when we wish to create curves.
# The in-context example for chair (ContExample in Table 4)
This is the corresponding program to create a chair backrest with an arc using the arc command (A) as well as a circular hole with the CUT and R command.
¡SOL¿
L: (3,0)
L: (3,5)
A: (0,5,120,1)
L: (0,0)
¡CUT¿
R: (1.5,2,1)
E: (-1,0,0,0,0,0,-0.5,NewBody,OneSided)
Remember the L:(0,0) is to close the loop, and that the non-cut sketch loop is drawn counterclockwise. Also note here that we made the cut radius equal to 1 to keep the cut within the scope of the original sketch. The angle 120 in the ’A: ’ command denotes the angle formed by the arc in degrees, and the 1 indicates the counterclockwise flag direction (SAME AS THE COUNTERCLOCKWISE SKETCH DIRECTION!).
Note that in this example the backrest was extruded in the X-axis which points outwards, to define the thickness of the backrest. This is different from the natural extrusion axis direction of the chair leg which is typically the z-axis to capture different chair heights.
# The reminder of the basic rules (Reminder in Table 4)
Now, using this language, produce a sequence of commands to form the following 3D rendering of the model. Keep the following three rules in mind: 1. It is VERY IMPORTANT to detect and reconstruct ALL PARTS of the object in the given image. So please DO NOT MISS ANY PARTS. Please pay attention and be careful. 2. COMMENT PARTS with DISTINCT LABELS in the code. If a part is repeated (like a leg), please give DISTINCT COMMENTS WITH NUMBERS (LIKE #LEG 1, #LEG 2, ETC.). 3. Give the FULL PROGRAM for ALL PARTS. EXPLICITLY WRITE OUT the program line by line. Do NOT say or use words like ”REPEAT”, ”FOLLOWING THE PATTERN”, ELLIPSIS ”…”. It is IMPORTANT TO WRITE OUT THE FULL PROGRAM. 4. Follow the coordinate system convention as given when creating the shapes, and PROVIDE AN EXPLANATION FOR THE EXTRUSION AXIS FOR EACH PART IN RELATION TO THE COORDINATE SYSTEM CONVENTION. Moreover, give its thickness. 5. Provide a description of the shape of the sketch in the comments.
# Chain-of-Thought prompting by asking GPT to enumerate all parts then write the program (CoT in Table 4)
Can you first enumerate all the parts that you see in the chair in the given image and then give me the corresponding FULL program for it?
A.1.3. Implementation Details of Img2CAD and TrAssembler
The embedding size for all $q$, $k$, and $v$ in our transformers is 256. Each transformer comprises a layer normalization, an 8-head self-attention layer, and three fully connected layers with a dimension of 256. For the part transformer encoder, we additionally concatenate the positional encoding with the input command tokens to distinguish the order of different commands within each part. We utilize the pretrained CLIP model with the ViT-B/32 architecture to encode the text label and the input image.
For multiple extrusions in the same part, e.g., two extrusions within a cut backrest (one for the outer sketch and the other for the inner cut), we concatenate them into the same block under the same part label. For different parts with the same label, such as two legs, we further sort these parts using their ground-truth bounding box center and concatenate a positional embedding after sorting. This ensures that the global transformer has sufficient information to distinguish between different legs and outputs different attribute parameters for each.
We use a batch size of 64 and a learning rate of $3\times 10^{-4}$ with a weight decay of $1\times 10^{-6}$. For each category, we train TrAssembler for 300 epochs.
A.1.4. Detailed Implementation of the Baselines
DeepCAD and DeepCAD-End2End
For both DeepCAD and DeepCAD-End2End, the image encoder is instantiated similarly to our Img2CAD model. This encoder includes a CLIP embedding module with a ViT-B/32 architecture and a Multi-Layer Perceptron (MLP) with 6 layers, ReLU activation, and a hidden dimension of 256. To adapt the original point cloud-conditioned DeepCAD from Wu et al. (2021) to an image-conditioned model, we replace the point cloud encoder with the image encoder. The model is initially fine-tuned on our dataset using an auto-encoding objective. Subsequently, the program encoder and decoder are frozen, and the image encoder is optimized to minimize the L2 distance between the vectors embedded by the program encoder and the image encoder, given the program-image pairs as input. For DeepCAD-End2End, both the image encoder and program decoder are optimized by minimizing the cross-entropy loss on the commands and the quantized bins of the parameters.
A.1.5. Detailed Implementation of Our Ablations
For the quantized version of Img2CAD discussed in Table 2, we follow DeepCAD by quantizing the continuous parameter space into 256 bins. We modify the final output head of our model to be a classifier with an output dimension of 256, and the input encoding layer for the parameters to be a learnable embedding dictionary with a vocabulary size of 256. All other components are kept the same as in the original Img2CAD model.
A.1.6. Details in Computing the Metrics
For our main results, we adopt Chamfer Distance (CD), Part Segmentation Accuracy (Seg Acc), and Part Segmentation mean Intersection over Union (Seg mIoU) as the metrics. Specifically, for Chamfer Distance, we first reconstruct the 3D mesh using the model-predicted CAD program and uniformly sample 8,000 points on the surface of the mesh. Similarly, we sample 8,000 points on the surface of the ground truth mesh, reconstructed from the ground truth CAD program in our test set. The Chamfer Distance is then computed as the average distance between these two point clouds over all points.
For segmentation metrics, we use the part segmentation labels from ShapeNet. After obtaining the point cloud sampled from the reconstructed mesh, we assign a part label to each point in the point cloud based on the label of its nearest neighbor in the point cloud of the same ShapeNet model. Segmentation Accuracy (Seg Acc) and mean Intersection over Union (Seg mIoU) are then computed using a PointNet pretrained on the ShapeNet part segmentation task as the segmentation network.
A.1.7. More Details on Dataset
We present the detailed statistics of our dataset, including the number of data points, and the maximum and average number of lines and parts in the CAD program, in Table A1. Additionally, we plot the distribution of the number of lines and parts across all three datasets in Figure A1. As shown in the statistics, our dataset provides a detailed semantic description of common objects with a large number of semantic parts. Moreover, our dataset is more challenging, with significantly more lines in the CAD program compared to the previous CAD dataset ABC (Koch et al., 2019). We also follow the split of PartNet (Mo et al., 2019b) for our dataset, resulting in 862/164, 2691/552, and 252/53 training/testing data points for chairs, tables, and cabinets, respectively.
Table A1. Detailed statistics of our dataset. The total number of data points (Size) and the maximum/average number of lines and parts are exhibited.
Chair Table Cabinet
Size Max/Ave # Lines Max/Ave # Parts Size Max/Ave # Lines Max/Ave # Parts Size Max/Ave # Lines Max/Ave # Parts
1026 107/58.14 15/9.44 3243 99/48.44 15/8.06 305 110/62.94 15/10.28
Figure A1. The distribution of the number of lines (top) and the number of parts (bottom) in the CAD program of our dataset.
A.2. More Experiment Results
A.2.1. Detailed Results of DeepCAD-AE
We provide the auto-encoding fitting results of DeepCAD-AE and DeepCAD-AE fine-tuned on our dataset with all three metrics (CD, Seg Acc, and Seg mIoU) in Table A2. For these two models, the input is our CAD-ified 3D point cloud and the output is the same point cloud, as suggested by the term ”auto-encoding.” Fine-tuning on our dataset improves the performance of DeepCAD on the CAD representation of common objects, including chairs, tables, and cabinets, highlighting the value of our curated dataset. It illustrates that the pretrained DeepCAD, trained only on the ABC dataset (Koch et al., 2019), struggles to understand domestic common objects. However, it can achieve better performance when fine-tuned on our introduced CAD-ified data of everyday common objects, which supports the usability of our dataset to supplement the existing ABC dataset (Koch et al., 2019), which focuses on industrial parts.
Table A2. DeepCAD auto-encoder fitting results on the test split of our dataset. We report Chamfer Distance (CD), Part Segmentation Accuracy (Seg Acc) and Part Segmentation mIoU (Seg mIoU) on all three categories. Fine-tuning on our dataset consistently improves all three metrics.
Chair Table Cabinet
DeepCAD-AE 0.4015 34.05 30.58 0.4054 44.22 30.34 0.3799 55.86 43.96
DeepCAD-AE-finetuned 0.2101 74.96 64.46 0.2031 77.35 63.96 0.3109 65.86 52.91
A.2.2. More Ablations on Conditional Structure Factorization
We propose using a nearest neighbor approach to retrieve the semantic labels of each part at inference time. Specifically, at inference time, we parse the GPT-4V’s output and extract the semantics as the text comments GPT-4V produces at the beginning of the CAD program for each part. Empirically, however, the semantics provided by GPT-4V can be noisy. To stabilize the semantic embeddings, we replace the inferred semantics with the nearest neighbor from the set of semantics that appear in our training set. The distance between semantics is measured by the L2 distance between their CLIP embeddings. We refer to this approach as GPT w/ Retrieval.
Since our model is trained on the CLIP embeddings of the semantics, it also generalizes well without retrieval, which we refer to as GPT w/o Retrieval. As shown in Table A3, for both Img2CAD and its quantized version, the retrieval procedure generally improves performance, especially on the semantic segmentation metrics, such as Seg Acc and Seg mIoU. Moreover, our model without retrieval still outperforms the baselines, such as DeepCAD-End2End, by a significant margin. This is because we embed the semantic labels using a pretrained CLIP model, which can generalize to unseen texts appearing at inference time.
Table A3. Ablation studies on semantics acquisition in conditional structure factorization. Performing nearest neighbor retrieval for the GPT output leads to performance increase especially on the semantic segmentation metrics, for both Img2CAD and its quantized version.
Quantized Semantics Acquisition Chair Table Cabinet
- GPT w/ Retrieval 0.1685 81.79 73.32 0.1655 83.75 69.55 0.2327 66.32 57.33
- GPT w/o Retrieval 0.1866 78.27 68.41 0.1691 82.58 67.56 0.2291 65.06 55.30
$\checkmark$ GPT w/ Retrieval 0.1912 79.82 73.22 0.1881 81.77 67.40 0.3169 63.89 56.12
$\checkmark$ GPT w/o Retrieval 0.2116 74.55 64.14 0.2035 78.99 63.95 0.2894 67.94 54.38
A.3. Additional Qualitative Results
We further include additional qualitative results. Figure A2 shows additional examples for our ablation study, while Figure A3 shows additional real world examples on the Pix3D dataset.
Figure A2. Additional qualitative comparisons of ablation studies on ShapeNet CAD-ified test set.
Figure A3. Additional Image to CAD results on Pix3D dataset.
A.4. Dataset Visualization
We include visualizations of some examples of our CAD dataset of common everyday objects as shown in Figure A4,  A5,  A6.
Figure A4. Dataset visualizations for chairs. For each pair, left: ShapeNet model, right: our CAD-ified model.
Figure A5. Dataset visualizations for tables. For each pair, left: ShapeNet model, right: our CAD-ified model.
Figure A6. Dataset visualizations for cabinets. For each pair, left: ShapeNet model, right: our CAD-ified model.
A.5. Limitations and Future Work
While CAD models are often used for mechanical parts, we found that VLMs perform less effectively on images of such objects. We intend to address this issue using in-context learning and LLM/VLM fine-tuning in future work. Additionally, GPT-4V sometimes makes mistakes and fails to provide the exact number of parts seen in the image, as shown in Figure A7.
Figure A7. Failure cases. GPT-4V can sometimes hallucinate or miss some structural parts. For each pair, left: input image; right: our model prediction.

### v2

@@ 7. Applications
7.1. Image-to-CAD in the Wild
Although our model is trained exclusively on synthetic ShapeNet CAD-ified models, it demonstrates the ability to generalize to real-world images. This is achieved through the foundational knowledge provided by VLMs. Figures 12 and 14 present qualitative results using images from Pix3D (Sun et al., 2018) and Google furniture images. This task is exceptionally challenging due to the difficulties associated with generalization and CAD representation complexity.
Figure 12. Image to CAD results on Pix3D dataset.:
Figure 13. Failure cases. Llama3.2 can sometimes hallucinate or miss some structural parts, and TrAssembler can struggle on predicting the accurate position on complex long-tailed examples (right-most).:
Figure 14. Image to CAD results on Google furniture images.:
Figure 15. Image conditioned CAD editing example. Users can easily create multiple variations of a table from a single-view image.:
7.2. Structure-Aware Editing.
A key advantage of our approach is its support for structure-aware editing in real-world scenarios, enabled by our interpretable CAD representation. As illustrated in Figure 15, users can intuitively modify an object’s CAD structure within an image using natural language instructions. The process begins by retrieving the CAD program using Img2CAD. This program is then provided to GPT-4o, which identifies the necessary modifications to achieve the user-specified outcome. It supports both structural changes, e.g., transforming a rectangular table into a round one, and continuous attribute adjustments, such as modifying the table length. GPT-4o exhibits strong reasoning abilities by understanding part relationships: it adjusts both the tabletop and the legs accordingly.
@@ 8. Conclusion
We propose a novel approach for reverse engineering 3D CAD models from input images by conditionally factorizing this challenging task into two sub-problems. First, we leverage VLMs (i.e., Llama3.2) to infer the discrete global structure with semantic information. We then introduce TrAssembler, which, conditioned on this discrete global structure, predicts the continuous attribute values of the CAD program. Additionally, we constructed a CAD dataset of common objects from ShapeNet to support the training of the continuous attribute prediction network. Our experiments demonstrate promising results, significantly outperforming baselines. Our method also makes a significant step towards generating CAD models from images in the wild, showcasing its potential for real-world applications.
9. Limitations and Future Work.
Although flow matching achieves strong results, the sampling time remains dependent on the number of inference timesteps. For instance, using the Euler ODE solver with 32 steps requires approximately 5 seconds to generate a single sample, highlighting a limitation of diffusion-like models. We anticipate that incorporating recent one-step diffusion methods could significantly reduce inference time, which we leave for future work.
Furthermore, perfect symmetry and connectivity are not guaranteed by our inference-time symmetry guidance. This guidance acts as a soft constraint, as each flow-matching step alternates between a standard flow update (which remains on the learned manifold) and a symmetry-score update (which encourages symmetry). This interleaving biases the CAD parameters toward symmetry while maintaining validity, resulting in improved symmetry and connectivity in most cases (Figure 8, Table 2), if not always. In future work, we plan to investigate methods for enforcing symmetry and connectivity directly within the flow-matching steps.
Despite finetuning, Llama3.2 remains imperfect. As shown in Figure 14, the predicted structure occasionally contains missing or hallucinated parts, indicating challenges in handling complex or long-tailed cases. To quantify the potential upper bound, we conducted an experiment where the predicted structure from Llama3.2 was replaced with the ground-truth CAD structure. This substitution led to a reduction in Chamfer Distance (CD) from 0.1174 to 0.1032. Future work could focus on enhancing the robustness of language models in predicting accurate and complete CAD programs.
Acknowledgements.
@@
This work is supported by ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a gift from the Flexiv corporation. Yang You also acknowledges the support from the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University.